Find the extreme values of f(x,y)=x^2+y on the surface x^2+y^2=1


1
Expert's answer
2021-06-02T11:01:58-0400

Solve equations ∇f= λ ∇g and g(x,y)=1 using Lagrange multipliers Constraint:


"g(x, y)=x^2+y^2=1"

Using Lagrange multipliers,


"f_x=\lambda g_x, f_y=\lambda g_y, g(x, y)=1"

which become


"2x=2x\lambda"

"1=2y\lambda"

"x^2+y^2=1"

If "x=0," then "y^2=1=>y=\pm1."

If "\lambda=1," then "y=\dfrac{1}{2}," and "x=\pm\dfrac{\sqrt{3}}{2}"

Therefore "f" has possible extreme values at the points "(0, -1), (0, 1),( -\dfrac{\sqrt{3}}{2}, \dfrac{1}{2})," and "(\dfrac{\sqrt{3}}{2}, \dfrac{1}{2})." Evaluating "f"at these four points, we find that 


"f(0, -1)=-1"

"f(0, 1)=1"

"f( -\dfrac{\sqrt{3}}{2}, \dfrac{1}{2})=\dfrac{5}{4}"

"f( \dfrac{\sqrt{3}}{2}, \dfrac{1}{2})=\dfrac{5}{4}"

Therefore the maximum value of "f" on the circle "x^2+y^2=1" is


"f( \pm\dfrac{\sqrt{3}}{2}, \dfrac{1}{2})=\dfrac{5}{4},"

the minimum value of "f" on the circle "x^2+y^2=1" is


"f( 0,-1)=-1"